# On the semiclassical bounce

### v1

@@ Introduction
Eq. (2) can be used to find a solution to the effective action, and its covariant nonlocal form [20] (see also further developments in [21] and [22]) is
@@
where $E_{4}=R_{\mu\nu\alpha\beta}^{2}-4R_{\alpha\beta}^{2}+R^{2}$ is the Gauss-Bonnet invariant, $G(x,y)$ is the conformally covariant Green function of the operator $\Delta_{4}\,=\,\Box^{2}+2R^{\mu\nu}\nabla_{\mu}\nabla_{\nu}-\frac{2}{3}\,R\Box+\frac{1}{3}\,R_{;\mu}\,\nabla^{\mu}$ and $\int_{x}=\int d^{4}x\sqrt{-g(x)}$.
@@
[…]
@@ Acknowledgments
W.C.S. is grateful to CAPES for supporting his Ph.D. project. The work of I.Sh. is partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico - CNPq under the grant 303635/2018-5.

### v2

@@ Introduction
Eq. (2) can be used to find a solution to the effective action, and its covariant nonlocal form [20, 21] (see also further developments in [22] and [23]) is
@@
where $E_{4}=R_{\mu\nu\alpha\beta}^{2}-4R_{\alpha\beta}^{2}+R^{2}$ is the Gauss-Bonnet invariant, $G(x,y)$ is the conformally covariant Green function of the operator $\Delta_{4}\,=\,\Box^{2}+2R^{\mu\nu}\nabla_{\mu}\nabla_{\nu}-\frac{2}{3}\,R\Box+\frac{1}{3}\,R_{;\mu}\,\nabla^{\mu}$ and $\int_{x}=\int d^{4}x\sqrt{-g(x)}$. It is possible to formulate the induced action in the covariant local form [20], including with two auxiliary scalar fields [24]. The last is the most useful formulation for many applications, such as classification of vacuum states [25] or the reaction of gravitational waves to the presence of higher derivatives [26]. A qualitatively similar representation, with certain simplifications [22, 23], should be most useful for the analysis of cosmological perturbations. We leave this part for the future work and, in the rest of this paper, will restrict the consideration by the basic elements of the cosmological model, i.e., the dynamics of the homogeneous and isotropic Universe. In this case, one can use a much simpler form of induced action, which is equivalent to (4) for this special metric.
@@
$\frac{\ddot{a}}{a}+\frac{\dot{a}^{2}}{a^{2}}\,=\,\frac{\mathcal{M}}{2a^{4}}+\frac{16\pi}{3M_{P}^{2}}\,\rho_{\Lambda}.$ (6)
@@
Another open question is the stability of the bounce model under discussion under the density and metric perturbations. This issue is typically complicated in all bounce models. The reason is that, in the vicinity of the bounce, the time derivative $\dot{H}$ is necessary positive and this leads to the violation of the Null Energy Condition (NEC). This feature may lead to instabilities in cosmic perturbations [38, 6] (see also [39] for an alternative discussion). In general, the analysis of the cosmological perturbations in a cosmological models is a necessary element of its development and this is especially true for models with a bounce [7]. Only the analysis of perturbations may show whether the given model is viable or possesses inconsistencies.
In the existing literature, there are strong indications of that the violation of NEC by quantum corrections may not lead to the inconsistencies [40] and that the same is true in the theories with scalar fields [41, 42]. Both arguments apply to our case. It is important to note that the perturbations should be analyzed not on the basis of a simplest non-covariant form of induced action (5), but using the covariant form (4), in the local representation. In this case, the theory always includes two auxiliary scalar fields [24, 25, 26] and, therefore, there are chances to arrive at the consistent model of bounce, including the perturbations free of pathologies, according to the criterion of [7]. We hope to come back to the detailed consideration of this issue and, as a first step, construct a new simplified formulation of the induced action with auxiliary fields, in a close future.
@@ Acknowledgments
Authors are very grateful to Patrick Peter for the interest to this work and very useful explanations concerning the role of cosmic perturbations in bounce models. W.C.S. wish to acknowledge CAPES for supporting his Ph.D. project. The work of I.Sh. is partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico - CNPq under the grant 303635/2018-5.